Aqueous hydrochloric acid reacts with solid sodium hydroxide to produce aqueous sodium chloride and liquid water . What is the theoretical yield of water formed from the reaction of 33.2 g of hydrochloric acid and 13.1 g of sodium hydroxide? Round your answer to 3 significant figures.
1
Expert's answer
2019-04-08T07:57:45-0400

"NaOH+HCl \to NaCl+H2O"

"n(NaOH)=\dfrac{13.1g}{40g/mol}=0.3275mol"

"n(HCl)=\dfrac{33.2g}{36.46g/mol}=0.9106mol"

NaOH is then the limiting reagent.

Theoretical yield"(H2O)=\dfrac{13.1g*18g/mol}{40g/mol}=5.90g"

Answer: theoretical yield(H2O) = 5.90g